# SWAP-less Implementation of Quantum Algorithms

### v1

@@ Acknowledgements
The authors thank P. Aumann, M. Lanthaler and G. B. Mbeng for valuable comments and discussions. This study was supported by the Austrian Research Promotion Agency (FFG Project No. FO999909249, FFG Basisprogramm). This research was funded in part by the Austrian Science Fund (FWF) under Grant-DOI 10.55776/F71 and Grant-DOI 10.55776/Y1067. This project was funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733. For the purpose of open access, the author has applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission. We acknowledge the support of the Federal Ministry of Education and Research (BMBF) within the framework of the program “Quantum technologies - from basic research to market”, QSOLID, under Grant No. 13N16154.

### v2

@@ Acknowledgements
The authors thank P. Aumann, M. Lanthaler and G. B. Mbeng for valuable comments and discussions. This study was supported by the Austrian Research Promotion Agency (FFG Project No. FO999909249, FFG Basisprogramm). This research was funded in part by the Austrian Science Fund (FWF) under Grant-DOI 10.55776/F71 and Grant-DOI 10.55776/Y1067. This project was funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733. For the purpose of open access, the author has applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission. We acknowledge the support of the Federal Ministry of Research, Technology, and Space (BMFTR) within the framework of the program “Quantum technologies - from basic research to market”, QSOLID, under Grant No. 13N16154.